OpenStudy (anonymous):

Solve on the interval [0, 2pi) sin3x=0

OpenStudy (anonymous):

\(\sin(3x)=0\) gives a few choices for \(3x\) for example since \(\sin(0)=0\) then \(3x=0\) or \(x=0\) is one solution

OpenStudy (anonymous):

another is \(\sin(\pi)=0\) and so \(3x=\pi\) making \(x=\frac{\pi}{3}\)

OpenStudy (anonymous):

there are others for example \(\sin(2\pi)=0\) making \(3x=2\pi\) ...

OpenStudy (anonymous):

Ah, I understand it now. I can't believe I didn't get it at first. Thank you for helping out!

OpenStudy (anonymous):

yeah i figured once you saw it you would see that it is easy. first solve for \(3x\) and then for \(x\) often the \(ax\) throws people off yw